atable with flat top and mass 10 kg is kept on a fl;at horizontal ground. a block of mass m=2 kg rests on the top of the table. a force F is applied at an angle of 30 degree and the block moves with an acceleration of root3 m/s.
1)find force exerted by block on eart
3)find force exerted by table on block
2)find force exerted by ground on table.
4)find net force acting on block

Grade:10

1 Answers

vikas askiitian expert
13 years ago

horizontal component force will provide accleration ...

Fcos30 = 31/2F/2  = FH = force in horizontal direction

ma = 31/2F/2

m=2kg , a = 31/2 so

F = 4N

net force acting in verticle direction = mg+Fsin30 = 22N

 

normal force bw table & block = 22N

 

normal force bw ground & table = 22+weight of table

                                            =22+100 = 122N

 

net force acting on block = [ N2 + (FH)2 ]1/2 

N is normal force bw block & table , FH is horizontal comp of applied force ....

                                   = [12 + 222]1/2 = 22.3N